The average public High School gas 468 students with a standard deviation of 87:


a. If a public school is selected, what is the probability that the number of students enrolled is greater than 400?


b. If a random sample of 38 public elementary schools is selected, what is the probability that the number of students enrolled is between 445 and 485?




1
Expert's answer
2021-03-31T16:42:57-0400

(a) The number of students enrolled in a randomly selected public school is a random variable "X", that has a normal probability distribution with parameters "\mu=468" and "\sigma=87". The value X=400 corresponds to z-value of a standard normal distribution (400-468)/87=-0.78. Then "P(X\geq 400)=1-\Phi(z)=\Phi(-z)=\Phi(0.78)=0.7823"

where "\Phi(x)=\frac{1}{\sqrt{2\pi}}\int\limits_{-\infty}^{x}e^{-t^2/2}dt" is cumulative distribution function of standard normal distribution.

(b) If a random sample of n=38 public elementary schools is selected, then the average number of student enrolled would be a normally distributed random variable "Y" with the mean "\mu=468" and the standard deviation "\sigma/\sqrt{n}=87/\sqrt{38}=14.11".

The value Y=445 corresponds to z-value of a standard normal distribution (445-468)/14.11=-1.63

The value Y=485 corresponds to z-value of a standard normal distribution (485-468)/14.11=1.20

Then

"P(445\leq Y\leq 485)=\Phi(1.20)-\Phi(-1.63)=0.8849-0.0516=0.8333"

Answer. (a) "P(X\geq 400)=0.7823";

(b) "P(445\leq Y\leq 485)=0.8333".